OpenStudy (maddieayeeee):

**Will Medal** Jennifer Kent receives 24 paychecks each year. Each check is for $863.50. If this is the amount of her check before deductions, what is Jennifer's annual salary?

alones (alones):

You have to solve it out?

OpenStudy (will.h):

Annual salary = all checks Annual salary = 24 * 863.50 Annual salary = ?

OpenStudy (maddieayeeee):

20,724?

OpenStudy (will.h):

indeed

OpenStudy (maddieayeeee):

Thanks. Also what is Jennifer Kent's weekly salary?

OpenStudy (maddieayeeee):

How do i find it?

OpenStudy (will.h):

how many weeks are there in 1 year?

OpenStudy (maddieayeeee):

52 weeks

OpenStudy (will.h):

great! so you will have to divide the annual salary by 52 weeks Weekly salary = 20,724/52 Weekly salary = ?

OpenStudy (maddieayeeee):

398.53

OpenStudy (will.h):

yes!

OpenStudy (maddieayeeee):

Thanks!

OpenStudy (will.h):

yw